Operations Report – January 4th

Crew 272 Operations Report 4-1-2023

SOL: 3

Name of person filing report: Mason Kuhn

Non-nominal systems: Suit Charger #1, Hab toilet,

Notes on non-nominal systems: I swapped the battery of suit 1 to suit 2, and then charged suit 2 (battery 1) with charger 1. The voltage reading after about 2 hours was unchanged at 10.38 V. I then moved the battery of suit 1 back to suit 1. I took charger 11 (since nobody seems to be using the lightweight suit) and attached it to suit 1 (battery 1). After 2 hours the voltage reading of battery 1 was 13.37 V. From this, I believe the issue is with charger 1 and that there is no issue with suit 1.

No new issues to report with the Hab toilet.


Spirit rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: Yes

Hours: 218.9

Beginning charge: 100%

Ending charge: 59%

Currently charging: Yes

Perseverance rover used: Yes

Hours: 253.4

Beginning charge: 100%

Ending charge: 51%

Currently charging: Yes

General notes on rovers: Curiosity and Perseverance were used on the long EVA today. On our way out

Summary of Hab operations: Before the toilet tank was emptied, crewmates noticed bad-smelling liquid coming from below the toilet. I emptied the toilet, and then that liquid was cleaned and sanitized extra. This was a result of waiting a bit too long to empty the toilet.


Water (static tank): 331

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: All suits except suit 11 are plugged in. Charger 11 is being used for suit 1 because charger 1 was determined to be weak. All other suits are charging. Suits 7-10 were used on the EVA from 1030 to 1340 (returned 50 minutes early). All radios 1-7 are fully charged and have been removed from the charger.

Summary of GreenHab operations: Watered and monitored all GreenHab crops, no new plantings. Monitored experimental crops growing in paper substrate, set up new experiments using potting soil and mixture of substrate gathered on today’s EVA with bleached and unbleached seeds.

WATER USE: 7 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 1g basil, 1g thyme, 1g oregano

Summary of ScienceDome operations:

Dual split: On

Summary of RAM operations: Retrieved a DMM and measuring tape around 0750 which have been in the lower deck of the Hab all day. I was checking suit charges and transmitter 9V battery charges throughout the day. The measuring tape has been used to track our water usage throughout the day (to find which operations use the most water) due to very high usage yesterday.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support:

  • Request for info on rover battery:
  • On today’s EVA, we noticed some strange battery data. On the way to our destination, both Perseverance and Curiosity were going around 80% speed on mode M (always in mode M), and the battery of each drained to 60-66% charge. We had to stop due to a rover reaching 60% charge and continue on foot. On our way back, both rovers were going around 60% speed on mode M (always in mode M), and the battery of each drained less than 10%. Going both directions, we slowed down significantly when the road terrain required slower speeds.

In short, is it normal for the rover battery to drain ~40% one way and ~10% on the way back? We were curious about if a slight speed decrease usually makes such a huge difference in battery usage.


Mason Kuhn

Crew Engineer

Copyright © The Mars Society. All rights reserved. | Main Site